14 (of 20) 4 O You did nor 10.00 points Problem 17-17 Aftertax comparison of preferred stock and other investments [LO17-5 has some excess cash that it would lke to invest in Treasury bonds at a 11 percent yield, (b) corporate bonds at a 14 percent yield: rate on dividends is 10 percent or (c) preferred stock at an 12 percent yielid Omega Corporation is in a 40 percent tax bracket and the tas 8401% 24 AM ^E de ENG 1/12/2019 O Type here to
a-1. Compute the aftertax yields for the three investment options. (Do not round intermediate calculations. Aftertax yields b. Corporate bonds c. Preferred stock 8401% 7.201% a-2. Which one of the three investments should she select based on the aftertax yields? O Preferred stock O Treasury bonds O Corporate bond References eBook & Resources Worksheet Problem 17-17 Aftertax comparison of preferred stock and other investments Learning Objective: 17-05 Preferred stock is an intermediate type of security that falls sor n debt and Type here to search
Answer #1

1)

After tax treasury bond yield = 0.11 ( 1 - 0.4) = 0.066 or 6.6%

After tax corporate bond yield = 0.14 ( 1 - 0.4) = 0.084 or 8.40%

After tax preferred stock yield = 0.12 ( 1 - 0.4) = 0.072 or 7.20%

2)

Corporate bond as it has the highest yield.